### v1

# Addressing Data Heterogeneity in Federated Learning with Adaptive Normalization-Free Feature Recalibration

Abstract
Federated learning is a decentralized collaborative training paradigm that preserves stakeholders’ data ownership while improving performance and generalization. However, statistical heterogeneity among client datasets poses a fundamental challenge by degrading system performance. To address this issue, we propose Adaptive Normalization-free Feature Recalibration (ANFR), an architecture-level approach that combines weight standardization and channel attention. Weight standardization normalizes the weights of layers instead of activations. This is less susceptible to mismatched client statistics and inconsistent averaging, thereby more robust under heterogeneity. Channel attention produces learnable scaling factors for feature maps, suppressing those that are inconsistent between clients due to heterogeneity. We demonstrate that combining these techniques boosts model performance beyond their individual contributions, by enhancing class selectivity and optimizing channel attention weight distribution. ANFR operates independently of the aggregation method and is effective in both global and personalized federated learning settings, with minimal computational overhead. Furthermore, when training with differential privacy, ANFR achieves an appealing balance between privacy and utility, enabling strong privacy guarantees without sacrificing performance. By integrating weight standardization and channel attention in the backbone model, ANFR offers a novel and versatile approach to the challenge of statistical heterogeneity. We demonstrate through extensive experiments that ANFR consistently outperforms established baselines across various aggregation methods, datasets, and heterogeneity conditions.
1 Introduction
Federated learning (FL) (McMahan et al., 2017) is a decentralized training paradigm which enables clients to jointly develop a global model without sharing private data. By preserving data privacy and ownership, FL holds significant promise for applications across various domains, including healthcare, finance, and mobile devices. However, a fundamental challenge in FL is statistically heterogeneous, i.e. non-independent and identically distributed (non-IID) client datasets. This heterogeneity can degrade the performance of the global model and slow down convergence (Li et al., 2020b; Hsu et al., 2019). Addressing it is thus critical for the success of FL in real-world scenarios.
Most prior relevant research has focused on developing novel aggregation algorithms that either aim to reduce client drift (Karimireddy et al., 2020) or tailor specific layers of the model to individual clients to accommodate data variations (Zhang et al., 2023). While these algorithmic strategies have advanced the field, they often overlook the impact of the underlying model architecture on performance under heterogeneity. Of particular interest is the role of normalization layers. Batch Normalization (BN) (Ioffe & Szegedy, 2015), although effective in centralized settings, has been shown to negatively affect performance in heterogeneous FL due to mismatched client-specific statistics and inconsistent parameter averaging (Wang et al., 2023; Guerraoui et al., 2024). In response, the use of alternative feature normalization methods like Group Normalization (GN) (Wu & He, 2018) and Layer Normalization (LN) (Ba et al., 2016) has been frequent (Hsieh et al., 2020; Reddi et al., 2021; Wang et al., 2021; Du et al., 2022). While these alternatives do not rely on mini-batch statistics, they offer inconsistent benefits, often resulting in slower convergence and lower performance compared to BN (Chen & Chao, 2021; Tenison et al., 2023; Zhong et al., 2024). Previous works have not designed models specifically tailored to combat heterogeneity, leaving a gap in FL research.
We address this gap in the image domain by proposing Adaptive Normalization-Free Feature Recalibration (ANFR), an architecture-level approach designed to enhance robustness in FL under data heterogeneity. ANFR combines weight standardization (Qiao et al., 2020) with channel attention mechanisms (Hu et al., 2018) to directly tackle the challenges posed by non-IID data. Weight standardization normalizes the weights of convolutional layers instead of activations, eliminating the dependence on mini-batch statistics that are problematic in FL. This makes the model less susceptible to mismatched client statistics and inconsistent averaging during aggregation. Channel attention modules produce learnable scaling factors for feature maps, enabling the model to suppress features that are inconsistent across clients due to heterogeneity and to emphasize those that are consistent and relevant. By integrating channel attention with weight-standardized models, ANFR enhances the model’s ability to focus on shared, informative features across clients. This combination boosts performance beyond the individual contributions of weight standardization and channel attention, enhancing class selectivity and optimizing the distribution of channel attention weights. ANFR operates independently of the aggregation method and is effective in both global and personalized FL settings, with minimal computational overhead. Furthermore, when training with differential privacy, ANFR achieves an appealing balance between privacy and utility, enabling strong privacy guarantees without sacrificing performance.
We validate the effectiveness of ANFR through extensive experiments on a diverse set of datasets, including medical imaging and natural image classification tasks, under various types of data heterogeneity. Our results demonstrate that ANFR consistently outperforms established baselines across different aggregation methods, datasets, and heterogeneity conditions. By focusing on architectural components, our approach is orthogonal to existing algorithmic advancements and represents a crucially underdeveloped area in FL research. The proposed ANFR model offers a robust and flexible solution to the challenge of statistical heterogeneity, contributing to the advancement of federated learning by enhancing performance, stability, and privacy-preserving capabilities.
2 Related Work
Since McMahan et al. (2017) introduced FL, most research has focused on developing novel aggregation algorithms to address challenges like data heterogeneity. In global FL (GFL), methods such as proximal regularization (Li et al., 2020a) and cross-client variance reduction (Karimireddy et al., 2020) aim to reduce client drift. Techniques like discouraging dimensional collapse through correlation matrix norm regularization (Shi et al., 2023), adopting relaxed adversarial training (Zhu et al., 2023), and performing amplitude normalization in frequency space (Jiang et al., 2022) have also been proposed. Recent promising ideas include constructing global pseudo-data to de-bias local classifiers and features (Guo et al., 2023), introducing concept drift-aware adaptive optimization (Panchal et al., 2023), and using hyperbolic graph manifold regularizers (An et al., 2023). In personalized FL (pFL), personalizing specific layers of the model can mitigate heterogeneity. The simplest approach shares all model parameters except the classification head (Arivazhagan et al., 2019). More advanced methods replace lower layers and mix higher ones (Zhang et al., 2023) or adjust mixing ratios based on convergence rate approximations (Jiang et al., 2024). While these algorithmic approaches have advanced both GFL and pFL, they often overlook the impact of the underlying architecture on performance.
Our work addresses this gap by examining how specific model components can enhance FL performance. This approach is orthogonal to algorithmic advancements, representing a crucially underdeveloped area. Previously, Qu et al. (2022) demonstrated performance gains by switching from a pre-trained convolutional network to a pre-trained vision transformer. Studies by Pieri et al. (2023) and Siomos et al. (2024) evaluated multiple architectures under different aggregation methods, showing that changing the backbone model can improve performance more than altering the aggregation method. However, these works did not design models specifically tailored to combat heterogeneity. In contrast, our method directly addresses data heterogeneity by integrating into the model architectural components that enhance robustness across diverse client distributions.
The normalization layer has been a focal point of component examination because Batch Normalization (BN) (Ioffe & Szegedy, 2015), the most widely used normalization technique, has been shown both theoretically (Li et al., 2021; Wang et al., 2023) and empirically (Hsieh et al., 2020; Du et al., 2022; Guerraoui et al., 2024) to negatively impact performance in heterogeneous FL settings. Mismatched local distributions lead to averaged batch statistics and parameters that fail to accurately represent any source distribution. Two primary approaches address this issue: modifying the aggregation rule for the BN layer or replacing it entirely. Some methods keep BN parameters local to each client (Li et al., 2021; Andreux et al., 2020) or stop sharing them after a certain round (Zhong et al., 2024). Others replace batch-specific statistics with shared running statistics when normalizing batch inputs to match local statistical parameters (Guerraoui et al., 2024) or leverage layer-wise aggregation to also match associated gradients (Wang et al., 2023). These methods still rely on decently sized batches to accurately approximate statistics and are incompatible with differential privacy. To replace BN, Group Normalization (GN) (Wu & He, 2018) has been frequently used (Hsieh et al., 2020; Reddi et al., 2021; Wang et al., 2021) since it does not rely on mini-batch statistics. Du et al. (2022) showed that Layer Normalization (LN) (Ba et al., 2016) performs better than GN in some FL settings. However, studies have shown these alternative normalization techniques offer inconsistent benefits over BN, depending on factors like dataset characteristics and heterogeneity levels (Tenison et al., 2023; Chen & Chao, 2021; Zhong et al., 2024). Moreover, GN requires tuning the number of groups for optimal effectiveness.
We propose circumventing these issues by applying weight standardization (Qiao et al., 2020) to normalize the weights of the model instead of the activations. Inspired by Brock et al. (2021a), who showed that such Normalization-Free (NF) models can train stably and perform on par with those employing BN in centralized learning, we explore this concept in FL. Previously, Zhuang & Lyu (2024) proposed an aggregation method specific to NF models for multi-domain FL with small batch sizes. Similarly, Siomos et al. (2024) showed that NF-ResNets improve upon vanilla ResNets under different initialization schemes and aggregation methods, while Kang et al. (2024) proposed a personalized aggregation scheme that replaces each BN layer with weight normalization (Salimans & Kingma, 2016) followed by a learnable combination of BN and GN.
Our method adaptively recalibrates the resulting feature maps using channel attention modules, such as the Squeeze-and-Excitation block (Hu et al., 2018). By doing so, we enhance the model’s ability to focus on relevant features across clients, effectively addressing data heterogeneity. Zheng et al. (2022) previously explored channel attention for FL in a different context, proposing a personalized FL method where a modified channel attention block is kept personal to each client. Unlike previous methods limited to specific aggregation strategies or settings, our approach can complement any heterogeneity-focused aggregation method, is effective even with large batch sizes, and supports various attention modules. Appendix C summarizes the differences between ANFR and related work. By integrating weight standardization with channel attention, ANFR provides a robust and flexible solution to data heterogeneity in FL, overcoming limitations of activation normalization techniques and complementing various aggregation strategies.
3 Adaptive Normalization-Free Feature Recalibration
We consider a FL setting with clients, each owning a dataset of image-label pairs and optimizing a local objective , where is a loss function and the model parameters. The global objective is learning a model that minimizes the aggregate loss:
Heterogeneity among can degrade the global model performance and slow convergence (Kairouz et al., 2021). In this study, we are interested in modifying the backbone model to address it. As they are the most widely used family, and they perform better or on par with others (Pieri et al., 2023; Siomos et al., 2024), we focus specifically on convolutional neural networks (CNNs). Let represent a batch of image samples with channels and dimensions . For a convolutional layer with weights and a kernel size of , the outputs are given by:
In typical CNNs, the activations are then normalized:
where are learnable parameters, is an indexing vector and is the set of pixels over which are computed. BN computes statistics along the axes, LN along , and GN along separately for each of groups of channels.
Channel attention (CA) mechanisms, like the Squeeze-and-Excitation (SE) block (Hu et al., 2018), recalibrate feature responses by modeling inter-channel relationships. The channel descriptor is obtained via Global Average Pooling (GAP):
This descriptor is then non-linearly transformed to capture dependencies between channels; in SE blocks this is done via the learnable weights and , where is a dimensionality reduction ratio:
yielding per-channel scaling factors which are applied to the normalized activations .
In the presence of data heterogeneity, CA can suppress features sensitive to client-specific variations and emphasize consistent ones. In earlier layers, consists of responses to filters detecting low-level features like colors and edges, while in later layers it contains class-specific features (Zeiler & Fergus, 2014). For the sake of explaining how CA impacts heterogeneous FL, we virtually partition filters into two distinct groups: those eliciting consistent features () and inconsistent ones (). Figure 1 illustrates an example. Both clients have images of airplanes and cars; Client 1’s images have predominantly blue backgrounds, while Client 2’s images have different backgrounds. Under this feature shift, edge-detecting filters produce consistent responses across both clients, thus belonging to , whereas filters sensitive to specific colors like blue activate differently across clients, forming . While both activation types are informative locally, inconsistent activations from cause conflicting gradients during FL training. This motivates our use of CA in this context: during training, CA can assign higher weights to and lower weights to without prior knowledge of which features belong to each set. The resulting adaptive recalibration aligns feature representations across clients, reducing gradient divergence and improving global model performance.
While CA mitigates the locality of convolution by accessing the entire input via pooling (Hu et al., 2018), if the normalization of is ill-suited to heterogeneous FL, the input to (4) becomes distorted, leading to sub-optimal channel weights:
Activation normalization techniques suffer from this issue. BN is known to be problematic in heterogeneous settings for two reasons: mismatched client-specific statistical parameters lead to gradient divergence—separate from that caused by heterogeneity—between global and local models (Wang et al., 2023); and biased running statistics are used at inference (Guerraoui et al., 2024). Both contribute to well-established performance degradation (Li et al., 2021; Du et al., 2022). Since and depend on batch-specific statistics, varies across clients due to local distribution differences, leading to inconsistent channel descriptors, which in turn results in non-ideal channel weights. Aside from data heterogeneity, BN needs sufficient batch sizes to estimate statistics accurately, and is incompatible with differential privacy; these are limiting factors in resource-constrained and private FL scenarios. GN and LN also have drawbacks: GN normalizes within fixed channel groups, which may not align with the natural grouping of features, limiting its effectiveness under heterogeneity. LN assumes similar contributions from all channels (Ba et al., 2016), which is generally untrue for CNNs, and clashes with our goal of reducing the influence of . Crucially, both normalize across channels to produce . This introduces additional channel inter-dependencies in (6), thus interfering with extracting representative channel descriptors.
To address these problems, we propose applying CA after normalizing the convolutional weights instead of the activations using Scaled Weight Standardization (SWS) from NF models (Brock et al., 2021a), which adds learnable affine parameters to weight standardization (Qiao et al., 2020):
Here, incorporates a learnable scale parameter and a fixed scalar depending on the networks’ non-linearity. We replace the normalized activation with . From (7) we observe that SWS does not introduce a mean shift (), and preserves variance () for the appropriate choice of , allowing stable training. By replacing normal convolutions with the ones described by (7), and following the signal propagation steps described in Brock et al. (2021a), we can train stable CNNs without activation normalization. We term this combination of weight standardization and channel attention Adaptive Normalization-Free feature Recalibration (ANFR). The input to (4) when using ANFR is:
Comparing (6) and (8), we note several advantages of ANFR. First, and are computed from convolutional weights, not the activations. Since weights are initialized identically and synchronized during FL, these weight-derived statistics are consistent across clients. Moreover, the second term of (8) now captures statistics of the input before convolution, providing an additional calibration point for CA and bypassing the effect of . By applying CA after SWS, we ensure channel descriptors are not distorted by batch-dependent statistics or cross-channel dependencies introduced by activation normalization. This allows CA to adjust channel responses effectively, improving the model’s capacity to learn stable feature representations that are consistent across clients with diverse data distributions. Therefore, the combination of SWS and CA overcomes the drawbacks of traditional normalization methods in federated learning, providing a novel and effective solution for improving model performance in the presence of data variability.
To further substantiate the effectiveness of integrating CA with weight standardization in addressing data heterogeneity, we conduct a mechanistic interpretability analysis comparing the effects of BN and SWS on class selectivity and attention weight variability. Specifically, we examine how well the ANFR model discriminates between classes both before111All networks are pre-trained on ImageNet. and after training on the heterogeneous ‘split-3’ partitioning of CIFAR-10 from Qu et al. (2022). This evaluation helps us understand how our method increases class discriminability under data heterogeneity. To isolate the effect of different components, we compare ANFR (using SWS with CA), to BN-ResNet (using BN), NF-ResNet (using SWS without CA) and SE-ResNet (using CA with BN). We quantify class selectivity using the class selectivity index (CSI) (Morcos et al., 2018), defined for each neuron as , where is the class-conditional activation that elicits the highest response and is the mean activation for all other classes. A right-skewed CSI distribution indicates higher class selectivity, crucial for effective classification under heterogeneous data. For models using CA (SE-ResNet and ANFR), we also examine attention weights distribution following Wang et al. (2020) to understand CA’s contribution to class discrimination.
Figure 2 shows CSI distributions for the last layer before the classifier, where class specificity is maximized in CNNs. Prior to FL training, incorporating CA in SE-ResNet with BN results in a small increase in class selectivity compared to BN-ResNet. When combining CA with SWS in ANFR, the change in class selectivity is negligible compared to NF-ResNet, indicating the minimal impact of CA at this stage. After training on heterogeneous data, we observe a notable shift: BN causes reduced class selectivity (compared to before training), evidenced by left-skewed distributions for BN-ResNet and SE-ResNet. Adding CA increases class selectivity for both normalization methods, but cannot fully mitigate BN’s negative effect due to inconsistently normalized inputs to the CA module (6). The ANFR model, however, shows a significant increase in class selectivity compared to NF-ResNet. This discrepancy manifests only after FL training, and indicates that combining CA and SWS in ANFR enhances the model’s ability to specialize in class discrimination under data heterogeneity.
For models using CA, we use the variability of attention weights across channels and classes as an indicator of adaptation: high variability suggests CA is actively re-weighing features to adapt to different class characteristics. Figure 3 shows that before FL training, both SE and ANFR models display high variability. This shows that CA provides a diverse and informative signal for both activation and weight normalization when heterogeneity is not a factor. However, after FL training, the attention weights of SE-ResNet become saturated, converging towards a value of 1 across all channels and classes. This saturation renders the attention mechanism inactive, turning it into an identity operation. Thus, the SE-ResNet model fails to preserve the discriminative power of CA under heterogeneity. Instead, the ANFR model maintains high variability in CA weights across channels and classes. This sustained variability implies that CA remains active and continues to provide class-discriminative signals when combined with weight standardization.
These mechanistic insights support our design choices. The adverse effects of BN under heterogeneous FL conditions are highlighted by the diminished class selectivity and inactive CA in SE-ResNet, while the ability of ANFR to maintain and improve class selectivity demonstrates that integrating CA with weight standardization effectively counters the impact of data heterogeneity. The enhanced class selectivity in ANFR correlates with improved downstream performance in heterogeneous FL settings, as we will show in Section 4. Additional details and extended CSI and attention weight results from other layers are presented in Appendix D.
4 Experiments
4.1 Experimental Settings
Datasets. We evaluate our approach on five classification datasets, including Fed-ISIC2019 (Ogier du Terrail et al., 2022) containing dermoscopy images from 6 centers across 8 classes where label distribution skew and heavy quantity skew is present; FedChest, a novel chest X-Ray multi-label dataset with 4 clients and 8 labels with label distribution skew and covariate shift; a partitioning of CIFAR-10 (Krizhevsky et al., 2009) which simulates heavy label distribution skew across 8 clients using the Kolmogorov-Smirnov (KS) ‘split-2’ as presented in Qu et al. (2022); CelebA (Liu et al., 2015) from the LEAF suite (Caldas et al., 2018) which is a binary classification task in a cross-device setting with a large number of clients, covariate shift and high quantity skew; and FedPathology, a colorectal cancer pathology slide dataset with 9 classes derived from Kather et al. (2019), featuring challenging concept drift as the images, which we do not color-normalize, were produced using two different staining protocols. FedChest contains images from PadChest (Bustos et al., 2020), CXR-14 (Wang et al., 2017) and CheXpert Irvin et al. (2019), which present one or more of 8 common disease labels. For FedPathology which we use for DP training in Section 4.3, we use Dirichlet distribution sampling (Hsu et al., 2019) with to simulate a moderate label distribution skew and partition the data among 3 clients. Each task covers a different aspect of the multi-faceted problem of data heterogeneity in FL, including different domains and sources of heterogeneity, to provide a robust test bed. Additional dataset details are presented in Appendix A.1, including instructions to replicate FedChest in B.1.
Compared models. We compare our approach with a typical ResNet (utilizing BN), a ResNet where BN is replaced by GN, a SE-ResNet (Hu et al., 2018), and a NF-ResNet. This selection isolates the effects of our architectural changes compared to using BN, using its popular substitution GN, and using weight standardization and CA separately. We choose a depth of 50 layers for all models to balance performance with computational expense. All models are pre-trained on ImageNet (Russakovsky et al., 2015) using the timm (Wightman, 2019) open-source library. ANFR follows the structure of NF-ResNet, with the addition of CA blocks in the same position as SE-ResNet. Except for Section 4.4, we employ Squeeze-and-Excitation (Hu et al., 2018) as the attention mechanism. Additional model details are provided in Appendix A.3.
Evaluated methods. We use 4 global FL (GFL) and 2 personalized FL (pFL) aggregation methods as axes of comparison for the models, each representing a different approach to model aggregation: the seminal FedAvg (McMahan et al., 2017) algorithm, FedProx (Li et al., 2020a), which adds a proximal loss term to mitigate drift between local and global weights, SCAFFOLD (Karimireddy et al., 2020), which corrects client drift by using control variates to steer local updates towards the global model, FedAdam (Reddi et al., 2021), which decouples server-side and client-side optimization and employs the Adam optimizer (Kingma & Ba, 2017) at the server for model aggregation, FedBN (Li et al., 2021) which accommodates data heterogeneity by allowing clients to maintain their personal batch statistics, and by construction is only applicable to models with BN layers, and FedPer (Arivazhagan et al., 2019) which personalizes the FL process by keeping the weights of the classifier head private to each client. We note our proposal is an architectural one which is aggregation method-agnostic, thus we selected these widely known aggregation methods to represent a spectrum of strategies, from standard averaging to methods addressing client drift and personalization. This provides a robust comparison concentrated on the model architectures.
Evaluation metrics. For Fed-ISIC2019, we report the average balanced accuracy due to heavy class-imbalance as in (Ogier du Terrail et al., 2022). For FedChest, a multi-label classification task with imbalanced classes, we report the mean AUROC on the held-out test set to account for varying class distributions, in addition to more extensive results in Appendix B.2. We report the average accuracy for the other 3 datasets. In pFL settings, the objective is providing good in-federation models. Consequently, we report the average metrics of the best local models, as suggested in (Zhang et al., 2023).
Implementation Details. We choose training hyper-parameters separately for each dataset by tuning the vanilla ResNet using the ranges detailed in Appendix A.2 and then re-using the same parameters for all models. In Fed-ISIC2019 clients use Adam with a learning rate of 5e-4 and a batch size of 64 to train for 80 rounds of 200 steps. In FedChest clients use Adam with a learning rate of 5e-4 and a batch size of 128 to train for 20 rounds of 200 steps. For DP-training in FedPathology, we set the probability of information leakage to , as is common, the noise multiplier to 1.1, the gradient max norm to 1.0, and train for 25 rounds, which is the point where the models have expended a privacy budget of . For CelebA and CIFAR-10 we follow the settings of Qu et al. (2022); Pieri et al. (2023) which were tuned by the authors. All experiments are run in a simulated FL environment with NVFLARE (Roth et al., 2022) and PyTorch (Paszke et al., 2019), using 2 NVIDIA A100 GPUs for training. We report the mean and standard deviation across 3 seeds.
4.2 Performance Analysis and Comparison
| Dataset | Method | Architecture | ||||
| BN-ResNet | GN-ResNet | SE-ResNet | NF-ResNet | ANFR (Ours) | ||
[…]
| Dataset | Method | Architecture | ||||
| BN-ResNet | GN-ResNet | SE-ResNet | NF-ResNet | ANFR (Ours) | ||
[…]
| FedBN | 82.820.06 | N/A | 81.840.28 | N/A | N/A | |
[…]
| FedBN | 83.380.12 | N/A | 83.330.14 | N/A | N/A | |
GFL scenario. Average results for all datasets, models, and GFL aggregation methods are presented in table 1. First, we observe that GN does not consistently outperform the vanilla ResNet, supporting our pursuit of a more reliable alternative. For instance, GN is outperformed by BN in half of the tested aggregation methods on Fed-ISIC2019 and FedChest. Second, the sub-optimality of CA operating on BN-normalized features is evident, as the SE model frequently performs worse than BN-ResNet, notably across all aggregation methods on FedChest. NF-ResNet shows strong performance across all tasks and methods, confirming the potential of replacing activation normalization with weight standardization in FL. However, our proposed ANFR model consistently outperforms NF-ResNet, often by a considerable margin. For example, on Fed-ISIC2019 with SCAFFOLD, ANFR surpasses NF-ResNet’s mean balanced accuracy by more than 3%. For the FedChest dataset, we employ a large batch size of 128 to maximize the probability that all classes are represented in each batch, following best practices. This choice incidentally biases the setting in favor of BN models since larger batches reduce inconsistent averaging of mini-batch statistics and BN parameters. Despite this bias, ANFR emerges as the top-performing model across aggregation methods. These results indicate that integrating CA with SWS networks provides significant performance gains, suggesting that channel attention is a crucial component in designing effective FL models.
pFL scenario. Table 2 presents the results for pFL scenarios on Fed-ISIC2019 and FedChest. In FedChest, where images are grayscale and we use a large batch size, FedBN and FedPer are virtually equal: BN-ResNet achieves an AUROC of 83.38% with FedBN and 83.39% with FedPer, indicating that the estimated BN statistics of the local models closely match the true ones. GN-ResNet attains 83.73% with FedPer, slightly outperforming BN-ResNet, but ANFR with FedPer is the most performant option across both aggregation methods, yielding a mean AUROC of 83.8%. Conversely, under the severe label and quantity skew on Fed-ISIC2019, employing FedBN improves performance over FedPer for models employing BN. ANFR achieves the highest balanced accuracy of 84.94% nonetheless. Notably, GN performs worse than BN on Fed-ISIC2019, and the ineffectiveness of combining BN and CA is further evidenced, as SE-ResNet is outperformed by BN-ResNet in all scenarios. These findings demonstrate that adopting ANFR enhances performance across both datasets, leading to the best overall models. Unlike the trade-offs observed with BN-FedBN and GN-FedPer combinations, ANFR consistently outperforms other architectures across varying levels of data heterogeneity.
Cross-device experiments on CelebA. Table 3 presents the results of our models on the cross-device setting of CelebA, which contains 200,288 samples across 9,343 clients. While the binary classification task is relatively straightforward for individual clients, it poses challenges at the server level due to the vast number of clients and significant quantity and class skews—some clients have only a few samples or labels from a single class. We observe that ANFR outperforms the baseline models, demonstrating its adaptability across diverse FL scenarios.
| Architecture | BN-ResNet | GN-ResNet | SE-ResNet | NF-ResNet | ANFR (Ours) |
| Average Accuracy | 82.21.21 | 85.410.68 | 85.550.84 | 88.170.3 | 88.910.28 |
4.3 Sample-level Differentially Private Training
In privacy-preserving scenarios involving differential privacy (DP) BN cannot be used because calculating mini-batch statistics violates privacy-preservation rules, and is customarily replaced by GN. We demonstrate the utility of ANFR compared to GN in such settings using the FedPathology set-up described in Section 4.1. We first train with strict sample-level privacy guarantees, employing a privacy budget of = 1, followed by training without privacy constraints ( = ) to illustrate the privacy/utility trade-off of each model.
| Privacy Budget | ||
| GN-ResNet | 84.792.72 | 67.275.08 |
| ANFR (Ours) | 84.473.08 | 81.110.33 |
From the results presented in Table 4, we observe that with an unrestricted privacy budget, GN and ANFR perform comparably. However, when a strict budget is enforced GN suffers a sharp performance decrease of 17%, as expected following previous research (Klause et al., 2022), whereas ANFR’s average accuracy is reduced by only 3%. ANFR’s robustness under DP may be attributed to its reliance on weight standardization, which has been shown to benefit from additional regularization (Brock et al., 2021b; Zhuang & Lyu, 2024) such as that provided by DP-SGD’s gradient clipping and gradient noising. Our experiments show DP training induces a regularization effect that disproportionately benefits NF models like ANFR, an observation also reported by De et al. (2022). These findings make ANFR a promising candidate for furthering development and adoption of DP training in FL, thereby enhancing the privacy of source data contributors, such as patients.
4.4 Attention Mechanism Comparison
Next, we investigate the impact of different attention mechanisms on performance. We compare the SE module used in previous sections with ECA (Wang et al., 2020), and CBAM (Woo et al., 2018). ECA replaces SE’s fully-connected layers with a more efficient 1-D convolution to capture local cross-channel interactions. CBAM combines channel and spatial attention and utilizes both max and average pooling to extract channel representations. From Table 5 we observe that even the lowest-performing module on each dataset outperforms all baseline models from Tables 1 and 3, proving the robustness of our approach. No single mechanism consistently performs best, making further exploration of attention module designs an interesting avenue for future work.
| CIFAR-10 | Fed-ISIC2019 | FedChest | CelebA | |
| NF-ResNet | 96.72 0.05 | 72.49 0.60 | 83.40 0.11 | 88.17 0.30 |
| ANFR (SE) | 97.42 0.01 | 74.78 0.16 | 83.49 0.14 | 88.91 0.28 |
| ANFR (ECA) | 97.13 0.11 | 75.07 0.48 | 83.62 0.10 | 89.07 0.43 |
| ANFR (CBAM) | 97.05 0.08 | 74.19 0.68 | 83.47 0.15 | 89.31 0.41 |
4.5 Qualitative Localization Performance Comparison
Finally, we assess the localization capability of each architecture after FL training with the best aggregation method on FedChest, SCAFFOLD. We compare the bounding box annotations provided by Wang et al. (2017) with Grad-CAM++ (Chattopadhay et al., 2018) heatmaps generated for samples labeled Atelectasis or Pneumonia from the FedChest test set. Figure 4 shows that ANFR’s heatmaps closely align with the annotated bounding boxes, indicating better localization of pathologies compared to baselines. This improved localization aids model interpretability, which is crucial in areas like medical imaging.
5 Conclusion
We introduce ANFR, a simple approach combining the strengths of weight standardization and channel attention to address the challenges of data heterogeneity and privacy at a design level in FL. ANFR fills a gap by being the first method to simultaneously work in GFL, pFL, and private FL scenarios while being compatible with any aggregation method and offering a robust increase in performance. Extensive experiments demonstrate the superior adaptability and performance of ANFR, as it consistently surpasses the performance of baseline architectures, regardless of the aggregation method employed. Our results position ANFR as a compelling backbone model suitable for both global and personalized FL scenarios where statistical heterogeneity and privacy guarantees are important concerns. Our findings highlight the need to look beyond aggregation methods as the core component of federated performance and the critical role of architectural innovations in reaching the next frontier in private and collaborative settings.
Reproducibility Statement
We have taken several steps to ensure reproducibility of our results. Details for the datasets we used and their pre-processing are in Appendix A.1. We also introduce a dataset of our own creation, FedChest, and provide detailed instructions on how to reproduce it in Appendix B. How we tuned our models is explained in Appendix A.2. This paper introduces a model, so we will provide ImageNet pre-trained weights for it upon acceptance. Details for all the models we use are in Appendix A.3. We will open-source all of our code upon acceptance.
Appendix A Additional Implementation Details
A.1 Datasets
Skin Lesion Classification on Fed-ISIC2019. Fed-ISIC2019 (Ogier du Terrail et al., 2022) contains 23,247 dermoscopy images from 6 centers across 8 classes and is a subset of the ISIC 2019 challenge dataset. We follow the original pre-processing, augmentation, loss, and evaluation metric of (Ogier du Terrail et al., 2022). This means the loss function is focal loss weighted by the local class percentages at each client, and the reported metric is balanced accuracy, as counter-measures against class imbalance. The augmentations used include random scaling, rotation, brightness changes, horizontal flips, shearing, random cropping to and Cutout (DeVries, 2017). We train for 80 rounds of 200 local steps with a batch size of 64. The clients locally use Adam (Kingma & Ba, 2017), a learning rate of 5e-4, and a cyclical learning rate scheduler (Smith, 2017). In terms of heterogeneity, Fed-ISIC2019 represents a difficult task due to class imbalance and heavy dataset size imbalance, with the biggest client owning more than 50% of the data and the smallest client 3%.
CIFAR-10. Krizhevsky et al. (2009) consists of 50,000 training and 10,000 testing images from 10 classes. We follow the setup of Pieri et al. (2023), specifically the ‘split-2’ partitioning where each client has access to four classes and does not receive samples from the remaining six classes. This means we train for 100 rounds of 1 local epoch with a batch size of 32. Clients use SGD with a learning rate of 0.03 and a cosine decay scheduler, in addition to gradient clipping to . During training the images are randomly cropped with the crop size ranging from 5% to 100% and are then resized to .
CelebA from LEAF. A partitioning of the original CelebA (Liu et al., 2015) dataset by the celebrity in the picture, this dataset contains 200,288 samples across 9,343 clients. The task is binary classification (smiling vs not smiling). We follow the setup presented in Pieri et al. (2023), training with 10 clients each round until all clients have trained for at least 30 rounds. The other settings are the same as those for CIFAR-10.
FedPathology Slide Classification Dataset. A colorectal cancer pathology slide dataset (Kather et al., 2019), consisting of training images of Whole Slide Image (WSI) patches with labels split among 9 classes, is used to simulate a federation of 3 clients. We mimic one of the most important challenges in the WSI field by not color-normalizing the images, which come from two different labs with differences in staining protocols. The original color-normalized validation set from Kather et al. (2019) is kept as a common validation set. We follow common practice (Hsu et al., 2019) to simulate label skew data heterogeneity by using a Dirichlet distribution with to partition the data. Since this artificial partitioning is random, we make sure to use the same seeds across architectures and privacy settings to compare on exactly the same partitioning instances. Our pipeline is built using Opacus (Yousefpour et al., 2022) and (, )-Renyi Differential Privacy (RDP) (Mironov, 2017). Following good practices, the probability of information leakage is set to where represents each client’s dataset size. The DP-specific hyper-parameters of the noise multiplier and gradient max norm are set to 1.1 and 1, respectively. Data augmentation includes random horizontal and vertical flips, random color jittering, and random pixel erasing. Clients use Adam with a learning rate of , training for 500 local steps with a batch size of 64. Federated training is stopped after 25 rounds, which is the point where both architectures have expended, on average, a privacy budget of . Finally, we train without using DP under the same settings to form a clearer picture of the privacy/utility trade-off of each model.
Chest X-Ray Multi-Label Classification on FedChest. Please refer to Appendix B.1.
A.2 Hyper-parameter Tuning
Hyper-parameters were optimized for the BN-ResNet and then the same parameters were used for all networks. The ranges were as follows:
Local Steps: {100, 200, 500}
Rounds: {20, 50, 75, 100}
Batch size: {32, 64, 128}
Gradient Clipping: {None, Norm Clipping to 1, Adaptive Gradient Clipping (Brock et al., 2021b)}
Learning rate: {}
Optimizer: {Adam, AdamW, SGD with momentum}
Scheduler: {None, OneCycleLR, Cosine Annealing, Cosine Annealing with Warm-up}
FedProx : {1e-3, 1e-2, 1e-, 2}
FedAdam Server learning rate: {5e-4, 1e-3, 1e-2, 1e-1}
Discussion. We found both FL aggregation methods that introduce hyper-parameters difficult to tune: FedProx Li et al. (2020a) made a negligible difference for small values and decreased performance as we increased it; the server learning rate in FedOpt has to be chosen carefully, as large (1e-2,1e-1) learning rates led to non-convergence and small ones to disappointing performance. Gradient clipping helped ANFR but was detrimental to the vanilla ResNet. We found the use of a scheduler to be very beneficial for performance, as well as making the optimizer and initial learning rate choice less impactful. We store the intermediate learning rate at each client between rounds and resume the scheduler, and also follow this for the momentum buffers of the adaptive optimizers.
A.3 Model Details
Table 6 presents pre-training details, parameter counts, multiply-accumulate counts (GMACs) and floating point operation counts (FLOPs) and ImageNet (Russakovsky et al., 2015) validation set top-1 performance for all of the compared models. For models which are pre-trained by us, links to the pre-trained weights will be made public after acceptance.
We note that the computational overhead of adding channel attention modules—which is already small—is offset by the abolishment of BN which is more expensive than SWS and incurs a bigger memory and time overhead (Brock et al., 2021b).
| Model | Parameters | GMACs | GFLOPs | IN-1K performance | Training Recipe |
| BN-ResNet-50 | 25.56 M | 4.09 | 8.21 | 78.81 | B |
| GN-ResNet-50 | 25.56 M | 4.09 | 8.24 | 80.06 | A1 |
| SE-ResNet-50 | 28.09 M | 4.09 | 8.22 | 80.26 | B |
| NF-ResNet-50 | 25.56 M | 4.09 | 8.32 | 80.22* | B |
| ANFR-50 (SE) | 28.09 M | 4.09 | 8.32 | 80.4 | B |
| ANFR-50 (ECA) | 25.56 M | 4.09 | 8.32 | 80.61 | B |
| ANFR-50 (CBAM) | 28.07 M | 4.1 | 8.33 | 80.37 | B |
Appendix B FedChest construction and additional results
B.1 Construction and hyper-parameters
To create FedChest we use three large-scale chest X-Ray multi-label datasets: CXR14 (Wang et al., 2017), PadChest (Bustos et al., 2020) and CheXpert (Irvin et al., 2019). To derive a common dataset format for all three, we need to take several pre-processing steps:
1.
We remove lateral views where present, keeping only AP/PA views.
2.
We discard samples which do not contain at least one of the common diseases, which are: Atelectasis, Cardiomegaly, Consolidation, Edema, Effusion, No Finding, Pneumonia, and Pneumothorax.
3.
We remove “duplicates” which, in this context, means samples from the same patient that have the same common labels but different non-common labels.
4.
We remove 5% from the edge of each image to avoid blown-out borders and artifacts.
5.
We resize the images to 224x224 pixels.
6.
We apply contrast-limited histogram equalization (CLAHE) to the images.
In addition to these common steps, some dataset-specific additional pre-processing steps are necessary, namely setting NaN and ‘uncertain’ labels of CheXpert to (not present), removing corrupted NA rows from CXR14, and removing corrupted images from PadChest.
After pre-processing, CheXpert has twice as many samples as the other datasets, so we further split it into two clients, cxp_young and cxp_old using the median age of the patient population (63 years), leading to a total of 4 clients with train/val/test splits of (given in thousands): for CXR14, for PadChest, for cxp_old and . The task is multi-label classification across the 8 common classes.
After tuning, clients perform 20 rounds of 200 local steps with a batch size of 128, the loss function is weighted Binary Cross-Entropy (BCE), and the optimizer Adam with a learning rate of , annealed over training. Data augmentation includes random shifts along both axes, random scaling and rotation, Cutout, and random cropping.
B.2 Additional Results
Further to the results presented in the main text, since some of the diseases have an unbalanced label distribution, and to also gauge model performance in deployment, we use the validation Receiver Operating Curve (ROC) to find the optimal class thresholds for each client using Youden’s Index (Youden, 1950). Having fixed the thresholds to these values, at test-time we measure the average accuracy and F1 score of each model and present the results in Table 7.
| Model | BN-ResNet-50 | GN-ResNet-50 | SE-ResNet-50 | NF-ResNet-50 | ANFR | |||||
| Metric | Accuracy | F1 | Accuracy | F1 | Accuracy | F1 | Accuracy | F1 | Accuracy | F1 |
[…]
| FedBN | 75.62 | 43.22 | N/A | N/A | 75.43 | 43.12 | N/A | N/A | N/A | N/A |
Appendix C Tabular Comparison with Related Work
Appendix D Extended CSI and Attention Weight Analysis
D.1 Setup details and performance
FL training is performed on the extremely heterogeneous ‘split-3’ partitioning of CIFAR-10 from Qu et al. (2022), which consists of 5 clients who each have samples only from 2 classes. The training parameters are the same as in Qu et al. (2022) and Section 4.1. All the compared models are pre-trained on ImageNet and have a depth of 50 layers, which results in 16 attention blocks for each model that uses channel attention. To calculate the channel attention weights and class selectivity index distributions, we use the entire test set of CIFAR-10, passing each class separately through the models to extract class-conditional activations; this is done both before and after FL training.
For channel attention weights, this allows us to store the distributions of weights of each model for each class and channel index. For the CSI, we query the nearest ReLU-activated feature maps before and after each channel attention block—or the equivalent points for the models that do not use such blocks. In timm (Wightman, 2019) terminology, we are referring to the output of act2 as before, and act3 as after. Comparing before and after distributions for the same network, allows us to isolate the effect of CA in the case of SE-ResNet and ANFR, and observe the baseline effect of moving through the convolutional block on the CSI distribution in BN-ResNet and NF-ResNet. Finally, the histogram of CSI values for each layer is used to draw an approximation of the continuous probability density function for the layer.
D.2 CSI plots of all layers
From Figure 5, which shows the CSI plots for every layer in the models, we make several observations regarding the class selectivity of each model.
SE-ResNet. Before FL training, CA reduces selectivity in all but the last block, in which it normalizes it. This is how CA was designed to function in the centralized setting, aiding feature learning in the first layers and balancing specificity and generalizability in the last layer (Hu et al., 2018). After FL training, the CSI distribution is much more left-skewed in the final block, showcasing how BN, under FL data heterogeneity, prohibits the network’s last layers from specializing compared to centralized training.
NF-ResNet. Before FL training we see that selectivity generally increases as we move towards the last layers. The CSI distribution of each layer after FL training is very similar to the one before it, indicating that replacing BN with SWS removes the limitation of the last layers to specialize.
ANFR. The distributions are generally similar to those of NF-ResNet except for some where CA reduces selectivity, adding to the evidence that part of the role of CA in centralized training is aiding general future learning. After heterogeneous FL training, ANFR inherits NF-ResNets robustness against heterogeneity, and by comparing the last layer of NF-ResNet and ANFR, we note that ANFR overall becomes more specialized.
D.3 Channel attention plots of all layers